OpenStudy (anonymous):

if f(x)=-6x+6, then f^-1(x)=

OpenStudy (saifoo.khan):

x = -6y + 6 Solve for y.

OpenStudy (anonymous):

i dont know how X(

OpenStudy (saifoo.khan):

Make "y" subject and you are done.

OpenStudy (anonymous):

i still dont understand what you are saying to do!

OpenStudy (anonymous):

i got 1 as the answer is that right

OpenStudy (anonymous):

basically reaarrange his given equation to solve for Y isolate Y to the side by itself

OpenStudy (anonymous):

subtract 6 from each side .........and then....

OpenStudy (anonymous):

you should get an equation not a value of 1

OpenStudy (anonymous):

x = -6y + 6 x - 6 = -6y -1/6x + 1 = y y = -1/6x + 1 is the answer